# Jet schemes, Newton polygons and continued fractions

### v2

@@ 4. Jet Schemes of Newton Plane Branches
In this section, we are interested in jet schemes of Newton non-degenerate plane branches. We begin by introducing combinatorial objects which are very important in this article: Newton polygons. A plane curve singularity $\mathcal{C}$ is the zero locus of an (reduced) element
@@
where $\mathcal{O}_{k^{2}}=k[x,y]$ has variables $x=x_{1}$ and $y=x_{2}$. One can associate with $\mathcal{C},$ more precisely with any $f\in\mathcal{O}_{k^{2}}$ its Newton polygon at the origin (notice that $f(0)=0$) as follows; let
@@
where Conv stands for the convex hull. The Newton polygon $\Gamma(f,x,y)$ is the union of the compact faces of (the boundary) of $\Gamma^{+}(f,x,y);$ it has one dimensional faces and vertices (00-dimensional faces). The dual fan of $\Gamma(f,x,y)$ (the dual Newton fan) is a fan (a collection of cones) whose support is $\mathbf{R}_{\geq 0}^{2}$ ($\mathbf{R}^{2}$ is the dual of the $\mathbf{R}^{2}$ that contains $\Gamma^{+}(f,x,y))$ and which is defined as follows: take $\omega\in\mathbf{R}^{2}_{\geq 0};$ and let
@@
$H_{\omega}(f)=\{(\alpha,\beta)\in\mathbf{R}^{2};\omega(\alpha,\beta)=\nu_{\omega}(f)\}$
With a face $\Phi$ of NP(f,x,y),(f,x,y), one associates the cone
$C_{\Phi}:=\{\omega\in\mathbf{R}^{2};\Phi\subset H_{\omega}(f)\};$
here $\Phi$ can be zero dimensional. The set $\{C_{\Phi},\Phi~\mbox{a face of NP}(f,x,y)\}$ is the dual Newton fan. The tropical variety $T(f)$ of $f$ is the union of the cones $C_{\Phi}$ associated to a compact segment $\Phi$ of the Newton Polygon of $f$. Let $p$ and $q$ be as in Section 3. A newton non-degenerate plane branch (or simply a Newton plane branch) $\mathcal{C}$ is the zero locus of an element
$f=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{ab}x^{a}y^{b}\in\mathcal{O}_{k^{2}}$
This coincides with the definition which we will see later in Section 5. In the case of $f$ defining a plane Newton non-degenerate branch, NP(f,x,y)(f,x,y) has two vertices $(q,0)$ and $(0,p)$ and one compact face which is the segment joining these two vertices. The dual of $(q,0)$ (respectively $(0,p))$ is the two dimensional cone generated by $(p,q)$ and $(0,1)$ (respectively $(1,0)$ and $(p,q))$. The dual cone of the compact segment is the one dimensional cone generated by $(p,q)$. This 1-dimensional cone is in fact $L_{(p,q)}$.
Set $\tau_{m+1}=\pi_{m+1,m}:\mathcal{C}_{m+1}\rightarrow\mathcal{C}_{m}$ the truncation morphism for $m\in\mathbf{N}_{0}$. Denote $irr(\mathcal{C}^{0}_{m})$ the set of irreducible components of the reduced scheme $\mathcal{C}^{0}_{m}$. Define the jet-component graph $Jet(f)$ of $f$ as follow:
@@
  • an edge of $Jet(f)$ joins two irreducible components $V_{m}\in irr(\mathcal{C}_{m}^{0})$ and $V_{m+1}\in irr(\mathcal{C}_{m+1}^{0})$ for some $m\in\mathbf{N}_{0}$, such that we have:
    $\tau_{m+1}(V_{m+1})\subseteq V_{m}$
An element of $irr(\mathcal{C}^{0}_{m})$ is called vertex of $Jet(f)$ at the level $m\in\mathbf{N}_{0}$.
@@
Let’s draw the jet components graph of the cusp ie $f=y^{2}-x^{3}$ in $\mathbf{A}^{2}$ (Figure 2) in two different ways. In the drawing on the right, introduce the notation $(\alpha,w)=\mathcal{H}^{\alpha}_{w}$. Note that we can have $w=\nu_{\alpha}(f)-1$ or $w=\infty$.
@@
The drawing on the left is obtained by the way that is described in the definition in the previous page. The one on the right is the new representation of the jet components graph that we’re introducing in this article. What makes the drawing on the right better than the one on the left are the following ideas:
@@
  • The drawing on the right is obtained by an algorithm whose steps are well-known. Moreover, it describes a link between the tropical variety and the irreducible components of the jet schemes (which isn’t the case for the drawing on the left).
  • The algorithm mentioned in the previous point is the same for any Newton non-degenerate plane branch (which we will see in the next example).
@@
  • When $k=\mathbf{C}$, the graph on the left is equivalent data with the embedded topological type of the singularity (defined by the cusp) as proven in [14]. Whereas the graph on the right gives richer algebraic data such as an exact description of the $\alpha$ (appearing in the irreducible components of the form $\mathcal{H}^{\alpha}_{m}$).
@@
  • We have to add to the graph on the left weights $(d,e)$ for each vertex in order to know which component vanishes at a certain level (that is when $d=e$) and which doesn’t. However, in the graph on the right, all nonvanishing components lie on the tropical variety.
@@
  • We can obtain the graph on the left from the one on the right: for every edge $\alpha-\alpha^{\prime}$ appearing in the graph on the right; where $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$ are the corresponding weights $w<w^{\prime}$, add on the edge $\alpha-\alpha^{\prime}$ an amount of $w^{\prime}-w-1$ dots corresponding to the $\mathcal{H}^{\alpha}_{m}$ where $w\leq m<w^{\prime}$.
@@
According to [14], every irreducible component $V_{m}\subset\mathcal{C}^{0}_{m}$ can be written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. In fact, using the notations in [14], if $\nu_{\alpha}(f)>m$ then $V_{m}=B_{m}$ and if $\alpha\in L_{(2,3)}$, $\nu_{\alpha}(f)\leq m$ then $V_{m}=\overline{Cont^{\alpha_{1}}(x)_{m}}$ doesn’t vanish and continues till $\infty$. This comes from: $\nu_{\alpha}(f)=6r$ when $\alpha=(2r,3r)$ : $r\in\mathbf{N}$.
@@
In the representation, we can simply replace $\mathcal{H}^{\alpha}_{m}$ by $\alpha$ and remove the text in red (for $f=y^{2}-x^{3}$) that is the name of an irreducible component that doesn’t vanish. This is because there is a correspondance between the irreducible components of $\mathcal{C}^{0}_{m}$ : $m\in\mathbf{N}$ and some weighted points in $\mathbf{R}_{\geq 0}^{2}$. Here is how it is defined: let $V_{m}\subset\mathcal{C}^{0}_{m}$ be an irreducible component written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. We associate to $V_{m}$ the weighted point $(\alpha,\infty)$ if $\alpha\in L_{(2,3)}$, $\nu_{\alpha}(f)\leq m$ and otherwise, we associate to $V_{m}$ the weighted point $\alpha,\nu_{\alpha}(f)-1]$.
@@
The weight of the point $\alpha$ is the level $w$ of the jet scheme $\mathcal{C}^{0}_{w+1}$ where the irreducible component $V_{w+1}$ vanishes: if $w<\infty$, $V_{w}=\mathcal{H}^{\alpha}_{w}$ is replaced by either $\mathcal{H}^{\alpha+(0,1)}_{w+1}$ [when $\alpha$ is below $L_{(2,3)}$] or $\mathcal{H}^{\alpha+(1,0)}_{w+1}$ [when $\alpha$ is above $L_{(2,3)}$]. If $w=\infty$, then $V_{m}$ doesn’t vanish.
We notice that if we unweight the points $(\alpha,w)\mapsto\alpha$ then these $\alpha$ follow the same pattern as in Figure 1. This concerns $f=y^{2}-x^{3}$.
@@
Let’s move on to any $f$ defining a Newton non-degenerate plane branch with $L_{(p,q)}$ as a tropical variety and $0<p<q$ coprime: let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Figure 3 is the staircase representation $J_{SC}(f)$ associated to $f$.
@@
Here is a description of $J_{SC}(f)$ in Figure 3: let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$ and recall the staircase notation $d_{1},...,d_{p}]]$ associated to $(p,q)$ from Theorem 1:
  • Start at $\lambda+(1,1)$;
  • For $1\leq i\leq p-1$, move $d_{i}$ vertical unit steps followed by one horizontal step;
  • Finally, move $d_{p}-1$ vertical steps; we reach $\lambda+(p,q)$;
  • Add a red arrow above $\lambda+(p,q)$;
  • Start over from $\lambda+(p,q)+(1,1)$.
Note that each time we move from a point $\alpha$ to $\alpha+(0,1)$ or $\alpha+(1,0)$, we delete $\alpha$ and we replace it by the new point. We weight each point $\alpha$ by $\nu_{\alpha}(f)-1$. The red arrows are never deleted and they are associated to points on the tropical variety with weight $\infty$. In the case where $0<q<p$, vertical steps are replaced by horizontal steps and vice versa.
The below proposition shows that $J_{SC}(f)$ determines $Jet(f)$.
@@
If $f,g$ define Newton plane branches with $J_{SC}(f)=J_{SC}(g)$ then we have $Jet(f)=Jet(g)$.
@@
Consider first an edge in Figure 3. Denote by $(\alpha,w)$ and $(\alpha^{\prime},w^{\prime})$ with $w<w^{\prime}<\infty$ the weighted vertices of this edge. Note that $\alpha^{\prime}=\alpha+(1,0)$ or $\alpha^{\prime}=\alpha+(0,1)$ with $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$. Add on the edge $\alpha\alpha^{\prime}$, $w^{\prime}-w-1$ vertices representing the big components defined by the same $\alpha$.
Now consider a nonvanishing component represented by $(\alpha,\infty)$. Add on the arrow above $\alpha$ a countably infinite amount of vertices. These represent the nonvanishing components for $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Finally, scale each weight $w$ by its level $m=w+1$.
The construction we just defined allows to obtain $Jet(h)$ from $J_{SC}(h)$ for every polynomial (or power series) $h$ which is non-degenerate (with respect to $x$ and $y$). Also, the construction doesn’t really depend on $h$ but on the weighted graph $J_{SC}(h)$. This allows to conclude that since $J_{SC}(f)=J_{SC}(g)$ then $Jet(f)=Jet(g)$ . ∎
@@ 5. Jet Schemes of Newton Plane Curves
such that for every face $\Phi$ of the Newton Polygon of $f$, if we set
@@
then the system of equations:
@@
has no solutions in the torus $(k^{\times})^{2}$.
In this study, it is crucial to know the description of Newton plane curves in order to compute their jet schemes over the origin of $k^{2}$. In this section, we will describe the equations of these curves.
Proposition 2 describes the equation of a Newton plane branch up to an isomorphism.
@@
Every curve contained in a Newton plane curve is also Newton non-degenerate. So, now, we describe the equations of Newton Plane curves through their defining polynomials (or power series) as well as the (equational) relations between their different branches.
@@
Newton Plane curves $\mathcal{C}\subset k^{2}$ are defined (up to an isomorphism) as such:
@@
    each branch $\mathcal{B}\subset\mathcal{C}$ is defined by a $g\in\mathcal{O}_{k^{2}}$ such that
    $g=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{a,b}x^{a}y^{b}$
    as in theorem 2.
@@
    Let $g^{\prime}=y^{p^{\prime}}-c^{\prime}x^{q^{\prime}}+\sum_{p^{\prime}a+q^{\prime}b>p^{\prime}q^{\prime}}c^{\prime}_{a,b}x^{a}y^{b}$ define another branch of $\mathcal{C}$. If $(p,q)=(p^{\prime},q^{\prime})$ then $c\neq c^{\prime}$.
In order to prove Theorem 2, we refer from [1] to an equivalent definition for Newton non-degenerate polynomials which is the following: let $f=\sum_{a,b}f_{a,b}\cdot x^{a}y^{b}\in k[x,y]$ be reduced and $w\in\mathbf{R}_{\geq 0}^{2}$, the initial form
@@
Equivalently, $f$ is Newton non-degenerate if and only if for all $(p,q)\in T(f)\cap\mathbf{N}^{2}$ coprime, $in_{(p,q)}(f)$ is the product of a non-zero constant by a monomial in $x,y$ (which could be equal to $1$) and a finite product of binomials in $x,y$: $\prod_{j}(y^{p}-c_{j}x^{q})$ where $j\mapsto c_{j}$ is injective.
Now, that we know the equations of the branches of a Newton plane curve and how they are related, let’s look at the irreducible components of $\mathcal{C}^{0}_{m}$ when $f=(y^{2}-x^{3})(y^{3}-x^{2})$ as seen in Figure 4. The tropical variety of $f$ is $T(f)=L_{(2,3)}\cup L_{(3,2)}$. Denote by $\sigma=L_{(2,3)}+L_{(3,2)}$ the cone such that $\partial\sigma=T(f)$.
@@
The special staircase walk for $f=(y^{2}-x^{3})(y^{3}-x^{2})$ determines the irreducible components of $\mathcal{C}^{0}_{m}$ : $m\in\mathbf{N}$.
@@
In fact, every irreducible component $V_{m}\subset\mathcal{C}^{0}_{m}$ is written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. We associate to $V_{m}$ then the weighted point $(\alpha,\infty)$ if $\alpha\in T(f)$, $\nu_{\alpha}(f)\leq m$ and otherwise, we associate to $V_{m}$ the weighted point $\alpha,\nu_{\alpha}(f)-1]$. Thus, we obtain Figure 4 by obeying to the following rules:
  • Start always at $(1,1)$;
  • If $\alpha\in\overset{\circ}{\sigma}$, remove $\alpha$ to replace it with $\alpha+(0,1)$ and $\alpha+(1,0)$;
  • If $\alpha\in L_{(2,3)}$, add a red arrow above $\alpha$ and add $\alpha+(1,0)$;
  • If $\alpha$ is above $L_{(2,3)}$ replace $\alpha$ by $\alpha+(1,0)$;
  • If $\alpha\in L_{(3,2)}$, add a red arrow above $\alpha$ and add $\alpha+(0,1)$;
  • If $\alpha$ is below $L_{(3,2)}$ replace $\alpha$ by $\alpha+(0,1)$.
The weighted point $(\alpha,w)$ represents the irreducible components $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ if $m<w<\infty$ that ”vanish” at the level $w+1$ and if $w=\infty$, then $V_{m}$ continue till $w=\infty$ without vanishing. These $V_{m}$, that do not vanish, are represented by a red arrow.
Now, let’s define the order of a power series $h=\sum_{k}h_{k}t^{k}\in k[[t]]\backslash\{0\}$. We set $ord_{t}(h)$ to be the least exponent $e$ such that $h_{e}\neq 0$. Otherwise if $h=0$, we set
$ord_{t}(0)=\infty$
@@
The semigroup of $f$ is the subsemigroup $\mathcal{S}(f)\subseteq\mathbf{N}^{r}_{0}$ whose elements are the:
@@
It is known [25] that the semigroup determines the embedded topological type of the singularity defined by $f$. When $f$ is Newton non-degenerate, we can rewrite this fact as:

### v3

@@ 4. Jet Schemes of Newton Plane Branches
In this section, we are interested in jet schemes of Newton non-degenerate plane branches. We begin by presenting a well-known combinatorial object: the Newton polygon. First, recall that a plane curve singularity $\mathcal{C}$ is the zero locus of an (reduced) element
@@
where $\mathcal{O}_{k^{2}}=k[x,y]$ has variables $x=x_{1}$ and $y=x_{2}$. More precisely, one can associate to any $f\in\mathcal{O}_{k^{2}}$ such that $f(0)=0$, its Newton polygon at the origin as follows. Let
@@
where Conv stands for the convex hull, then the Newton polygon $\Gamma(f,x,y)$ is the union of all compact faces of (the boundary of) $\Gamma^{+}(f,x,y);$ it has 1-dimensional faces (line segments) and 00-dimensional faces (vertices). The dual fan of $\Gamma(f,x,y)$ (the Newton dual fan) is a fan (a collection of cones) whose support is $\mathbf{R}_{\geq 0}^{2}$ and which is defined as follows: take $\omega\in\mathbf{R}^{2}_{\geq 0};$ and let
@@
$H_{\omega}(f)=\{(\alpha,\beta)\in\mathbf{R}^{2}\omega(\alpha,\beta)=\nu_{\omega}(f)\}.$
With a face $\Phi$ of $\Gamma(f,x,y),$ one associates the cone
$C_{\Phi}:=\{\omega\in\mathbf{R}^{2};\Phi\subset H_{\omega}(f)\}.$
The set $\{C_{\Phi},\Phi~\mbox{a face of }\Gamma(f,x,y)\}$ is the Newton dual fan. The tropical variety $T(f)$ of $f$ is the union of the cones $C_{\Phi}$ associated to each compact segment $\Phi$ of the Newton Polygon of $f$. Let $p$ and $q$ be co-prime integers such that $0<p<q$ (as in Section 3). A Newton non-degenerate plane branch (or simply a Newton plane branch) $\mathcal{C}$ is the zero locus of an element
$f=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{ab}x^{a}y^{b}\in\mathcal{O}_{k^{2}}.$
(This definition coincides with the one defined later in Section 5). In the case of $f$ defining a plane Newton non-degenerate branch, $\Gamma(f,x,y)$ has two vertices $(q,0)$ and $(0,p)$ and one compact face which is the line-segment joining these two vertices. The dual of $(q,0)$ (respectively $(0,p))$ is the 2-dimensional cone generated by $(p,q)$ and $(0,1)$ (respectively $(1,0)$ and $(p,q))$. Also, the dual cone of the compact segment is the one dimensional cone generated by $(p,q)$. This 1-dimensional cone is in fact $L_{(p,q)}$.
Next, we are going to recall from [17], the definition of the jet-components graph of a plane curve singularity.
Set $\tau_{m+1}=\pi_{m+1,m}:\mathcal{C}_{m+1}\rightarrow\mathcal{C}_{m}$ the truncation morphism for $m\in\mathbf{N}_{0}$, and denote by $irr(\mathcal{C}^{0}_{m})$ to be the set of irreducible components of the reduced scheme $\mathcal{C}^{0}_{m}$. Then, the jet-component graph $Jet(f)$ of $f$ can be defined as follows:
@@
  • an edge of $Jet(f)$ joins two irreducible components $V_{m}\in irr(\mathcal{C}_{m}^{0})$ and $V_{m+1}\in irr(\mathcal{C}_{m+1}^{0})$ for some $m\in\mathbf{N}_{0}$ whenever $\tau_{m+1}(V_{m+1})\subseteq V_{m}.$
Finally, an element of $irr(\mathcal{C}^{0}_{m})$ is called a vertex of $Jet(f)$ at the level $m\in\mathbf{N}_{0}$.
It is worth noting that in $Jet(f)$ there is a redundancy of irreducible components of the form $\mathcal{H}^{\alpha}_{m}$. In fact, one may have $\mathcal{H}^{\alpha}_{n}$ and $\mathcal{H}^{\alpha}_{m}$ as components with $n\neq m$. This allows us to construct a new graph where some of these components are removed without losing any information. The next example illustrates these ideas.
@@
Let us draw the jet-components graph of the cusp i.e. $f=y^{2}-x^{3}$ in $\mathbf{A}^{2}$ in the two ways mentioned earlier (See Figure 2 below). First, we introduce the notation $(\alpha,w)=\mathcal{H}^{\alpha}_{w}$ and we recall that $w=\nu_{\alpha}(f)-1$.
@@
The drawing on the left is obtained by the first way and the one on the right is the new representation of the jet components graph. What makes the drawing on the right more appealing than the one on the left might be the following reasons:
@@
  • The drawing on the right is obtained by an algorithm whose steps are described in Section 7. It describes a link between the tropical variety and the irreducible components of the jet schemes (which isn’t the case for the drawing on the left). In addition, this algorithm is the same for any Newton non-degenerate plane branch (which we will see in the next example).
@@
  • When $k=\mathbf{C}$, the graph on the left determines the embedded topological type of the singularity (defined by the cusp in our case) as proven in [14]. Whereas the graph on the right gives richer algebraic data such as an exact description of the $\alpha$ (appearing in the irreducible components of the form $\mathcal{H}^{\alpha}_{m}$).
@@
  • The weights $(d,e)$ on the left should be added on each vertex in order to know which component vanishes at a certain level (that is when $d=e$) and which does not. However, in the graph on the right, all nonvanishing components lie on the tropical variety.
@@
  • The graph on the left can be obtained from the one on the right as follows. For every edge $\alpha\alpha^{\prime}$ appearing in the graph on the right (where $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$ are the corresponding weights with $w<w^{\prime}$), an amount of $w^{\prime}-w-1$ dots should be added on the edge $\alpha\alpha^{\prime}$. These dots correspond to $\mathcal{H}^{\alpha}_{m}$ for $w\leq m<w^{\prime}$.
@@
Now, we justify why both representations are essentially the same. According to [14], every irreducible component $V_{m}\subset\mathcal{C}^{0}_{m}$ can be written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for some unique $\alpha\in\mathbf{N}^{2}$. Using the same notation as in [14], we note that if $\nu_{\alpha}(f)>m$ then $V_{m}=B_{m}$ and if $\alpha\in L_{(2,3)}$ and $\nu_{\alpha}(f)\leq m$, then $V_{m}=\overline{Cont^{\alpha_{1}}(x)_{m}}$ does not vanish and continues to $\infty$. This is due to the fact that $\nu_{\alpha}(f)=6r$ when $\alpha=(2r,3r)$ for $r\in\mathbf{N}$.
@@
First, we replace $\mathcal{H}^{\alpha}_{m}$ by $\alpha$ and remove the text in red (for $f=y^{2}-x^{3}$) that is the name of an irreducible component that does not vanish. This is because there is a correspondance between the irreducible components of $\mathcal{C}^{0}_{m}$ : $m\in\mathbf{N}$ and some weighted points in $\mathbf{R}_{\geq 0}^{2}$. Here is how it is defined: let $V_{m}\subset\mathcal{C}^{0}_{m}$ be an irreducible component written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. We associate to $V_{m}$ the weighted point $(\alpha,\infty)$ if $\alpha\in L_{(2,3)}$, $\nu_{\alpha}(f)\leq m$ and otherwise, we associate to $V_{m}$ the weighted point $\alpha,\nu_{\alpha}(f)-1]$.
It’s worthy to notice that there are two kinds of irreducible components for $\mathcal{C}^{0}_{m}$ which can be seen in the next table.
@@
The weight of the point $\alpha$ is the level $w$ of the jet scheme $\mathcal{C}^{0}_{w+1}$ where the irreducible component $V_{w+1}$ vanishes: if $w<\infty$, $V_{w}=\mathcal{H}^{\alpha}_{w}$ is replaced by either $\mathcal{H}^{\alpha+(0,1)}_{w+1}$ (when $\alpha$ is below $L_{(2,3)}$) or $\mathcal{H}^{\alpha+(1,0)}_{w+1}$ (when $\alpha$ is above $L_{(2,3)}$). If $w=\infty$, then $V_{m}$ does not vanish.
Concerning $f=y^{2}-x^{3}$, we notice that if we remove the weights from all the points $(\alpha,w)\mapsto\alpha$ then these $\alpha$’s follow the same pattern as in Figure 1.
Next, we consider the case of a Newton plane branch having a semi-line as a tropical variety.
@@
Let $f$ define a Newton non-degenerate plane branch with $L_{(p,q)}$ as a tropical variety and $0<p<q$ coprime with the staircase notation $d_{1},...,d_{p}]]$ associated to $(p,q)$ from Theorem 1. An inspection shows that Figure 3 is the staircase representation $J_{SC}(f)$ associated to $f$ which can be described as follows. For $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$, then
  • the walker starts at $\lambda+(1,1)$,
  • for $1\leq i\leq p-1$, the walker moves $d_{i}$ unit steps vertically followed by one step horizontally,
  • the walker takes $d_{p}-1$ steps vertically until reaching $\lambda+(p,q)$,
  • next, we add a red arrow above $\lambda+(p,q)$,
  • finally, the walker moves to $\lambda+(p,q)+(1,1)$ and restarts over.
Note that each time the walker moves from a point $\alpha$ to either $\alpha+(0,1)$ or $\alpha+(1,0)$, then $\alpha$ is deleted and replaced by the new point. In addition, each point $\alpha$ has a weight of $\nu_{\alpha}(f)-1$. Also, the red arrows are never deleted and they are associated to points on the tropical variety with weight $\infty$. In the case where $0<q<p$, vertical steps are replaced by horizontal steps and vice versa.
@@
The next proposition shows that $J_{SC}(f)$ determines $Jet(f)$.
@@
If $f,g$ define some Newton plane branches with $J_{SC}(f)=J_{SC}(g)$ then we have $Jet(f)=Jet(g)$.
@@
Consider first an edge in Figure 3. Denote by $(\alpha,w)$ and $(\alpha^{\prime},w^{\prime})$ with $w<w^{\prime}<\infty$ the weighted vertices of this edge. Note that $\alpha^{\prime}=\alpha+(1,0)$ or $\alpha^{\prime}=\alpha+(0,1)$ with $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$. Add on the edge $\alpha\alpha^{\prime}$, $w^{\prime}-w-1$ vertices representing the big components defined by the same $\alpha$. Now consider a nonvanishing component represented by $(\alpha,\infty)$. Add on the arrow above $\alpha$ a countably infinite amount of vertices. These represent the nonvanishing components for $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Finally, scale each weight $w$ by its level $m=w+1$.
This construction that we just described allows us to obtain $Jet(h)$ from $J_{SC}(h)$ for every polynomial (or power series) $h$ which is non-degenerate (with respect to $x$ and $y$). One of the advantages of this construction is that it depends only on the weighted graph $J_{SC}(h)$ and not on the polynomial $h$. This allows us to conclude that since $J_{SC}(f)=J_{SC}(g)$ then $Jet(f)=Jet(g)$. ∎
@@ 5. Jet Schemes of Newton Plane Curves
where for every face $\Phi$ of the Newton Polygon of $f$, the restriction of $f$ to $\Phi$
@@
has no critical points in the algebraic torus $(k^{\times})^{2}$ i.e. the system of equations:
@@
has no solutions in this torus.
It is worth mentioning that it is crucial in our study to know the description of Newton plane curves in order to compute their jet schemes over the origin of $k^{2}$. In this section, we will describe the equations of these curves.
First, we start with Proposition 2 that describes the equation of a Newton plane branch up to an isomorphism.
@@
Recall that every curve which is contained in a Newton plane curve is also Newton non-degenerate. So we can describe the equations of Newton plane curves through their defining polynomials (or power series) as well as the (equational) relations between their different branches.
@@
Any Newton plane curve $\mathcal{C}\subset k^{2}$ is defined (up to an isomorphism) by:
@@
    each branch $\mathcal{B}\subset\mathcal{C}$ is given by some $g\in\mathcal{O}_{k^{2}}$ such that
    $g=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{a,b}x^{a}y^{b}.$
@@
    If $g^{\prime}=y^{p^{\prime}}-c^{\prime}x^{q^{\prime}}+\sum_{p^{\prime}a+q^{\prime}b>p^{\prime}q^{\prime}}c^{\prime}_{a,b}x^{a}y^{b}$ is another branch of $\mathcal{C}$ with $(p,q)=(p^{\prime},q^{\prime}),$ then $c\neq c^{\prime}$.
In order to prove Theorem 2, we recall from [1] an equivalent definition for Newton non-degenerate polynomials which can be described as follows. Let $f=\sum_{a,b}f_{a,b}\cdot x^{a}y^{b}\in k[x,y]$ be reduced and $w\in\mathbf{R}_{\geq 0}^{2}$, the initial form
@@
Equivalently, $f$ is Newton non-degenerate if and only if for all coprime $(p,q)\in T(f)\cap\mathbf{N}^{2}$, then $in_{(p,q)}(f)$ is the product of a non-zero constant by a monomial in $x,y$ (which might be equal to $1$) and a finite product of binomials in $x,y$: $\prod_{j}(y^{p}-c_{j}x^{q})$ where $j\mapsto c_{j}$ is injective.
Now, that we know the equations of the branches of a Newton plane curve and how they are related, let us consider the following example.
@@
The tropical variety of $f=(y^{2}-x^{3})(y^{3}-x^{2})$ is $T(f)=L_{(2,3)}\cup L_{(3,2)}$. Denote by $\sigma=L_{(2,3)}+L_{(3,2)}$ the cone whose boundary is $T(f)$. The special staircase walk for $f=(y^{2}-x^{3})(y^{3}-x^{2})$ determines the irreducible components of $\mathcal{C}^{0}_{m}$ for $m\in\mathbf{N}$.
@@
In fact, every irreducible component $V_{m}\subset\mathcal{C}^{0}_{m}$ is written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for some unique $\alpha\in\mathbf{N}^{2}$. Then we associate with $V_{m}$ the weighted point $(\alpha,\infty)$ if $\alpha\in T(f)$, $\nu_{\alpha}(f)\leq m$ and otherwise, we associate with $V_{m}$ the weighted point $\alpha,\nu_{\alpha}(f)-1]$. Thus, we obtain Figure 4 from the following rules:
  • The walker starts always at $(1,1)$,
  • if $\alpha\in\overset{\circ}{\sigma}$, then we replace $\alpha$ with $\alpha+(0,1)$ and $\alpha+(1,0)$,
  • if $\alpha\in L_{(2,3)}$, then we draw a red arrow above $\alpha$ and add $\alpha+(1,0)$,
  • if $\alpha$ is above $L_{(2,3)}$ then we replace $\alpha$ by $\alpha+(1,0)$,
  • if $\alpha\in L_{(3,2)}$, then we draw a red arrow above $\alpha$ and add $\alpha+(0,1)$,
  • if $\alpha$ is below $L_{(3,2)}$ then we replace $\alpha$ by $\alpha+(0,1)$.
The weighted point $(\alpha,w)$ represents the irreducible components $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ that ”vanish” at the level $w+1$ for $m<w<\infty$. For $w=\infty$, then $V_{m}$ continues to $w=\infty$ without vanishing. Those $V_{m}$ that do not vanish, are represented by a red arrow.
Next, we define the order of a power series $h=\sum_{k}h_{k}t^{k}\in k[[t]]\backslash\{0\}$. First, we set $ord_{t}(h)$ to be the least exponent $e$ such that $h_{e}\neq 0$. Otherwise if $h=0$, we set
$ord_{t}(0)=\infty.$
@@
The semigroup of $f$ is the subsemigroup $\mathcal{S}(f)\subseteq\mathbf{N}^{r}_{0}$ whose elements are given by:
@@
It is known [25] that the semigroup determines the embedded topological type of the singularity defined by $f$. When $f$ is Newton non-degenerate, we can rewrite this fact as follows.